In a cyclotron, the diameter of the pole faces is 100 cm and the magnetic field between the pole faces is 0.50 T. The cyclotron is used for accelerating protons. Calculate the kinetic energy of proton in eV and speed of the proton as it emerges from the cyclotron. Also determine the cyclotron frequency. 


1
Expert's answer
2022-05-15T17:03:30-0400

"W=\frac{B^2V}{2\mu_0}=\frac{\pi B^2d^3}{12\mu_0}=52~kJ."